Strong base is dissolved in 605 mL of 0.200 M weak acid (Ka=4.32×10−5) to make a buffer with a pH of 3.99. Assume that the volume remains constant when the base is added.

HA(aq)+OH−(aq)⟶H2O(l)+A−(aq)

Calculate the pKa value of the acid, and determine the number of moles of acid initially present.

pka =

initial amount (mol HA) =

A-/HA=

how many moles of strong base were initially added?

When the reaction is complete, what is the concentration ratio of conjugate base to acid?

[A−][HA]=[A−][HA]=

How many moles of strong base were initially added?
Answer #1

Strong base is dissolved in 605 mL of 0.200 M weak acid (Ka=4.32×10−5) to make a buffer with a pH of 3.99. Assume that the volume remains constant when the base is added.

HA(aq)+OH−(aq)⟶H2O(l)+A−(aq)

1)

Ka = 4.32 x 10^-5

pKa = -log Ka = -log (4.32 x 10^-5)

pKa = 4.36

2)

initial moles of HA = 605 x 0.200 / 1000

initial moles of HA   = 0.121

3)

pH = pKa + log [A- / HA]

3.99 = 4.36 + log [A- / HA]

log [A- / HA] = - 0.37

[A- / HA] = 0.427

4)

A- / 0.121 = 0.427

A- = 0.0516 mol

moles of strong base added = 0.0516